OpenStudy (ksaimouli):

2^x+2^x+2^x+2^x=2^7. What is the value of x

OpenStudy (anonymous):

4(2^x) = 2^7 = 128 2^x = 32 x = 5

OpenStudy (anonymous):

All good now, @ksaimouli ?

OpenStudy (ksaimouli):

Can't we use rules of expo

OpenStudy (anonymous):

ok, (4)(2^x) = 2^7 (2^2)(2^x) = 2^7 2^x = (2^7)/(2^2) 2^x = 2^(7-2) 2^x = 2^5 x = 5

OpenStudy (ksaimouli):

I got it but how did u multi 4

OpenStudy (ksaimouli):

4(2^x)

OpenStudy (anonymous):

Because there were 4 identical terms on the left.

OpenStudy (ksaimouli):

Thx got it

OpenStudy (anonymous):

2^x+2^x+2^x+2^x = (4)(2^x) ok, just got your message. uw!